A girl runs and then dives into a pool with an average speed of 5.6 m/s before hitting the water. Calculate the height from which she dove?
1
Expert's answer
2020-10-19T13:21:18-0400
"v^2=2gh\\5.6^2=2(9.8)h\\h=1.6\ m"